Calculate the mass of carbon dioxide gas produced by the complete combustion of 692 g of Isooctane in 1600 g of oxygen
1
Expert's answer
2020-05-15T09:37:20-0400

2C8H18 + 25O2 = 16CO2 + 18H2O

n(C8H18) = 692g/114g/mol = 6.07g

n(O2) = 1600g/32g/mol = 50 mol

C8H18 is the limiting reactant, O2 is in the excess.

n(CO2) = 6.07*16/2 = 48.56 mol

m(CO2) = 48.56mol*44g/mol = 2136.7 g